# FalconFS: Distributed File System for Large-Scale Deep Learning Pipeline

### v1

@@ 6.4 Impact of Client Memory Budget
Figure 13: Random file traversal in a large directory tree.
In this section, we evaluate the impact of client memory budget on DFS performance under typical DL training workloads — specifically, random file traversal in a large directory tree. We initialize an 8-level directory tree structure where each intermediate directory contains ten subdirectories and each last-level directory contains ten 64 KiB files. This configuration yields a total of 11.1 million directories and 100 million files. Each DFS runs four metadata servers and twelve storage nodes. Ten client nodes, each running a 256-thread client process, read all files in independent random orders.
@@
Figure 13 shows the throughput of each DFS under different memory budgets and Figure 13 presents the composition of metadata requests. Notably, while Lustre and FalconFS explicitly send open requests to open files, CephFS sends lookup requests for file open. For simplicity, we count CephFS’s lookup requests to files as open in Figure 13.
We make the following observations: First, the performance of stateful-client DFSs, including CephFS, Lustre and FalconFS-NoBypass, is sensitive to the client memory budget. There is a 1.4–1.5$\times$ performance gap between the 10% and 100% memory budget configurations. When the memory budget is constrained, fewer directories can be cached on the client-side, leading to more frequent lookups, which increase the number of requests per file access and degrade throughput.
Second, FalconFS achieves consistently high performance even under tight memory budget and outperforms FalconFS-NoBypass, demonstrating that stateless-client design effectively boosts performance. As shown in Figure 13, FalconFS generates constant number of requests to the metadata servers as cache size varies. Compared with FalconFS-NoBypass, FalconFS reduce the number of metadata requests by 22.7%–45.9%, and improves the throughput by 0.24–0.94$\times$. Compared with CephFS and Lustre, FalconFS improve the throughput by 2.92–4.72$\times$ and 2.08–3.34$\times$ respectively.
6.5 Impact of Burst IO
We evaluate the impact of directory bursts on DFS performance, a common access pattern in DL labeling workloads. We deploy four metadata servers and 12 storage nodes, and uses a single 256-thread client node to access pre-created 64 KiB files in bursts. A burst is defined as a sequence of accesses to files within the same directory, with adjacent bursts targeting different directories. Figure 14 presents the results.
@@ 6.5 Impact of Burst IO
We observe performance degradation in CephFS’s read and write operations, as well as Lustre’s read performance, as the burst size increases. This occurs because large bursts cause instantaneous load imbalance across the metadata servers. In contrast, FalconFS does not suffer from large bursts, as it evenly distributes the metadata of files within the same directory, achieving good scalability. JuiceFS’s performance also does not degrade with increasing burst size, as there is a constant load imbalance among its metadata engine nodes.
@@ 6.6 Load Balance in Real Workloads
Table 3 summarizes, for each workload, the number of files, the maximum and minimum ratios of inodes on metadata servers, and the number of exception entries utilized to achieve the distribution. As shown in the table, most workloads exhibit a small maxmin gap with zero exception entries, indicating that filename hashing alone is sufficient to achieve balanced inode distribution in these cases. This is because such workloads — typically datasets — have large directory size, which is friendly to filename hashing.
We further examine the three workloads that require exception entries. The “Linux 6.8 code tree” contains many files with identical names; however, applying path-walk redirection to the two most frequent filenames (i.e., “Makefile” and “Kconfig”) suffices to balance the distribution. These two filenames occur 2,945 and 1,690 times respectively, accounting for 5.55% of all files in total. In the FSL home traces, FalconFS achieves balanced load after applying path-walk redirection to the most frequent filename, which appears 8,112 times and accounts for 1.24% of all files.
@@ 6.7 Performance Analysis
The contribution of the stateless-client architecture to overall performance is demonstrated by comparing FalconFS with FalconFS-NoBypass in §6.4. In this experiment, we further analyze other design configurations by evaluating three setups: the full FalconFS, no inv, and no merge, incrementally reducing the design features. The no inv configuration disables invalidation-based synchronization, wrapping mkdir operations in a distributed transaction to atomically create all dentry replicas across all MNodes. The no merge setup disables concurrent request merging in addition to the changes in no inv, requiring worker threads to fetch and execute requests one a time. For this evaluation, we deploy four MNodes and use LibFS clients to saturate them, with each client accessing its own private directory. Figure 15 presents the peak throughput of the mkdir operation.
Compared to the full FalconFS, no inv decreases throughput by 86.9%, as the distributed transaction requires multiple rounds of broadcasts, incurring significant overhead. no merge reduces throughput by an additional 91.8%, as request dispatching from the connection pool to the worker threads becomes a bottleneck due to shared request queue contention.
@@
In most cases, hybrid metadata indexing enables one-hop operations. However, there are corner cases that require two hops: (a) operations on non-existent paths, (b) operations on path-walk redirected filenames, and (c) operations issued with a stale exception table. Figure 15 illustrates how these scenarios affect the performance of the getattr operation. Compared to the one-hop common case, these corner cases result in a 36.8%–49.6% decrease in performance due to the additional hop.
@@ 6.8 End-to-End Performance
In this section, we evaluate the end-to-end performance of FalconFS in DL workloads for both labeling and training tasks. For the labeling task, we replay a trace captured from a production environment. In this trace, labeling tasks read raw images from the DFS and write segmented images back to the DFS. Figure 16 shows the distribution of file sizes in the trace, and Figure 16 presents the completion time of the trace replay. Although we do not replay GPU computation, the measured completion time closely approximates the end-to-end runtime, as computation overlaps with I/O and I/O remains the bottleneck. Compared to other DFSs, FalconFS reduces the runtime by 23.8%–86.4%.
@@
Figure 17: Accelerator utilization for Resnet-50 model training.
For the training task, we evaluate DFS performance using the MLPerf Storage Benchmark [28]. The benchmark is configured to simulate training a ResNet-50 model on 10 million files distributed across 1 million directories, with each file sized at 112 KiB, using varying numbers of GPUs. The total dataset size is 1,064  GiB, and files are read using direct I/O. Figure 17 shows the accelerator utilization (AU) of each DFS as the number of GPUs increases. JuiceFS is omitted because its throughput drops to zero during dataset initialization. Taking 90% AU as the threshold, FalconFS supports up to 80 GPUs, while Lustre supports only 32 GPUs, and CephFS does not meet the threshold. With 80 to 128 GPUs, FalconFS achieves training throughput speedups of 11.09–11.81$\times$ over CephFS and 0.99–1.23$\times$ over Lustre.
@@ 7 Related Works
To accelerate path resolution, DFSs typically adpots client-side metadata caching [37, 36, 25, 48, 4, 15, 27]. InfiniFS [27] reduces the latency penalty of cache misses by resolving multiple path components in parallel; however, it cannot mitigate request amplification. Giraffa [39] and CalvinFS [41] locate file inodes by full path hashing, which makes directory renaming hard to implement. HDFS [38] performs path resolution on a centralized namenode and thus has scalability issues. HopsFS [33] performs path resolution at a proxy layer, which looks up all path components in parallel from a distributed database, but still suffers from constant request amplification. Our approach differs in that FalconFS addresses scalability, request amplification and metadata indexing issues through a client-stateless architecture and hybrid metadata indexing.
@@ 8 Conclusion
We propose FalconFS, a distributed file system with a client-stateless architecture for DL workloads. Evaluations show that FalconFS achieves up to 4.72$\times$ better throughput of small file random access and up to 11.81$\times$ higher GPU utilization in deep learning model training over CephFS and Lustre.

### v2

@@ 6.4 Impact of Client Memory Budget
Figure 14: Random file traversal in a large directory tree. The x-axis represents the ratio of client metadata cache size to the total size of all directories’ inodes and dentries.
In this section, we evaluate the impact of client memory budget on DFS performance under typical DL training workloads — specifically, random file traversal in a large directory tree. We initialize an 8-level directory tree structure where each intermediate directory contains ten subdirectories and each last-level directory contains ten 64 KiB files. This configuration yields a total of 11.1 million directories and 100 million files. Each DFS runs four metadata servers and twelve data nodes. Ten client nodes, each running a 256-thread client process, read all files in independent random orders.
@@
Fig. 14 shows the throughput of each DFS under different memory budgets, and Fig. 14 presents the composition of metadata requests. Notably, while Lustre and FalconFS explicitly send open requests to open files, CephFS sends lookup requests for file open. For simplicity, we count CephFS’s lookup requests to files as open in Fig. 14.
[…]
Notably, even with a 100% cache, FalconFS-NoBypass is still 19.4% slower than FalconFS. Although this cache is large enough to hold all directory inodes in the workload, file inodes contend for the cache. Consequently, directory lookups still frequently miss the cache and generate remote requests.
6.5 Impact of Transient Skewness
We evaluate the impact of transient skewness on DFS performance, a common access pattern in DL labeling workloads (§2.4). We deploy four metadata servers and twelve data nodes, and use a single 256-thread client node to access pre-created 64 KiB files in bursts. A burst is defined as a sequence of accesses to files within the same directory, with adjacent bursts targeting different directories. Fig. 15 presents the results.
@@ 6.5 Impact of Transient Skewness
We observe a degradation in the read and write performance of CephFS and Lustre, as the burst size increases. This occurs because large bursts cause instantaneous load imbalance across the metadata servers. In contrast, FalconFS does not suffer from large bursts, as it evenly distributes the metadata of files within the same directory, achieving good scalability. JuiceFS’s performance also does not degrade with increasing burst size, as there is a constant load imbalance among its metadata engine nodes.
@@ 6.6 Load Balance in Real Workloads
Tab. 3 summarizes, for each workload, the number of files, the maximum and minimum ratios of inodes on metadata servers, and the number of exception entries utilized to achieve the distribution. As shown in the table, most workloads exhibit a small maxmin gap with zero exception entries, indicating that filename hashing alone is sufficient to achieve balanced inode distribution in these cases. This is because such workloads — typically datasets — have a large directory size, which is friendly to filename hashing.
We further examine the three workloads that require exception entries. The “Linux 6.8 code tree” contains many files with identical names. However, applying path-walk redirection to the two most frequent filenames (i.e., “Makefile” and “Kconfig”) suffices to balance the distribution. These two filenames occur 2,945 and 1,690 times, respectively, accounting for 5.55% of all files in total. In the FSL home traces, FalconFS achieves balanced load after applying path-walk redirection to the most frequent filename, which appears 8,112 times and accounts for 1.24% of all files.
@@ 6.7 Performance Analysis
The contribution of the stateless-client architecture to overall performance is demonstrated by comparing FalconFS with FalconFS-NoBypass in §6.4. In this experiment, we further analyze other design configurations by evaluating three setups: the full FalconFS, no inv, and no merge, incrementally reducing the design features. The no inv configuration disables invalidation-based synchronization, wrapping mkdir operations in a distributed transaction to atomically create all dentry replicas across all MNodes. The no merge setup disables concurrent request merging in addition to the changes in no inv, requiring worker threads to fetch and execute requests one at a time. For this evaluation, we deploy four MNodes and use LibFS clients to saturate them, with each client accessing its own private directory. Fig. 16 presents the peak throughput of the mkdir operation.
Compared to the full FalconFS, no inv decreases throughput by 86.9%, as the distributed transaction requires multiple rounds of broadcasts, incurring significant overhead. no merge reduces throughput by an additional 91.8% due to increased per-request overhead.
@@
In most cases, hybrid metadata indexing enables one-hop operations. However, there are corner cases that require two hops: (a) operations on non-existent paths, (b) operations on path-walk redirected filenames, and (c) operations issued with a stale exception table. Fig. 16 illustrates how these scenarios affect the performance of the getattr operation. Compared to the one-hop common case, these corner cases result in a 36.8%–49.6% decrease in performance due to the additional hop.
@@ 6.8 End-to-End Performance
We evaluate the end-to-end performance of DFSs in DL workloads for both labeling and training tasks. Each DFS has four metadata servers and twelve data nodes.
The labeling task.
We replay a trace from Huawei’s labeling cluster. In this trace, labeling tasks read raw images from the DFS and write segmented images back to the DFS. Fig. 17 shows the distribution of file sizes in the trace, and Fig. 17 presents the runtime of the trace replay. Although we do not replay the computation, the replay runtime closely approximates the end-to-end runtime, as computation overlaps with I/O, and I/O is the bottleneck. Compared to other DFSs, FalconFS reduces the runtime by 23.8%–86.4%.
@@
(a) Accelerator utilization.
(b) Runtime breakdown.
Figure 18: Accelerator utilization and runtime breakdown for Resnet-50 model training. In Fig. 18, the stripes mark the computation time, and the bars above the stripes represent the time waiting for I/O.
The training task.
[…]
@@ 7 Related Works
To accelerate path resolution, DFSs typically adopt client-side metadata caching [38, 37, 26, 49, 4, 15, 28]. InfiniFS [28] reduces the cache misses penalty by resolving multiple path components in parallel; however, it cannot mitigate request amplification. Giraffa [40] and CalvinFS [42] locate inodes by full path hashing, which makes directory renaming hard to implement. HDFS [39] performs path resolution on a centralized namenode and thus has scalability issues. HopsFS [34] performs path resolution at a proxy layer, which looks up all path components in parallel from a distributed database, but still suffers from constant request amplification. Our approach differs in that FalconFS addresses scalability, request amplification, and metadata indexing issues through a client-stateless architecture and hybrid metadata indexing. Mantle [24], a concurrent work of FalconFS, adopts a similar client-stateless metadata service to object storage services. While FalconFS scales out path resolution to all NModes, Mantle resolves path in a per-namespace single node and has scalability issues.
@@ 8 Conclusion
[…]
Acknowledgements
We sincerely thank our shepherd, Ken Birman, and the anonymous reviewers for their constructive comments and insightful suggestions. This work is supported in part by the National Natural Science Foundation of China (No. 62132014), the Fundamental Research Funds for the Central Universities, and Huawei Technologies. Mingkai Dong (mingkaidong@sjtu.edu.cn) and Junbin Kang (kangjunbin1@huawei.com) are the corresponding authors.
@@ A.1 Theoretical Analysis
A.2 Orthogonal Task-Level Optimizations
Previous studies have proposed task-level optimizations that change the way in which data is shuffled to make the I/O pattern more friendly to the DFS [46, 21]. Specifically, they group data objects into partitions, and shuffle the order of partitions and the order of objects in each partition separately for each epoch, in order to reduce the scope of the random access footprint.
FalconFS’s optimization is orthogonal to these task-level optimizations. While the task-level optimizations require engineering efforts on the training framework layer to implement and constrain the way in which data are shuffled, FalconFS satisfies the data demand of training tasks through filesystem-level optimization, which is transparent to upper-layer tasks and leaves sufficient room for the tasks to conduct orthogonal designs and optimizations.